Solve the system of equations using elimination. x+y= -8, -x+4y= 13
Answer by oberobic(2304)

x +y = -8
-x + 4y =13
.
multiply the first equation by 4 and then subtract
.
4x +4y = -32
-x +4y = 13
--------------
5x +0y = -45
.
x = -9
.
-(-9) + 4y = 13
9 + 4y = 13
4y = 4
y = 1
.
check
x + y = ??
-9 +1 = -8
correct
.
check
-x +4y = ??
-(-9) +4(1) = 9+4 = 13
correct
.
Answer: x=-9, y=1
.
Done.